I need help with this Word Problem:
A particular style of shoe costs the dealer $81 per pair. At what price should the dealer mark them so he can sell them at a 10% discount off the selling price and still make a 25% profit?
Thanks for your help!

Answer by stanbon(75887)

A particular style of shoe costs the dealer $81 per pair. At what price should the dealer mark them so he can sell them at a 10% discount off the selling price and still make a 25% profit?
-------------
Let the selling price be "x".
EQUATION:
x + 0.10x = 81 + 0.25*81
1.1x = 1.25*81
x = $92.05 (required selling price)
=============
Cheers,
stan H.